OpenStudy (ledah):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6

OpenStudy (anonymous):

multiply the first eqn by 2 then addition will cancel out the y's

OpenStudy (anonymous):

6x-2y=6 -2x+2y=6 4x=12 x=12/4 x=3